OpenStudy (anonymous):

I need help with algebra 2 please! I will fan & medal :)

OpenStudy (anonymous):

easier if you post a question

OpenStudy (anonymous):

After a dreary day of rain, the sun peeks through the clouds and a rainbow forms. You notice the rainbow is the shape of a parabola. The equation for this parabola is y = -x2 + 36. Create a table of values for a linear function. A drone is in the distance, flying upward in a straight line. It intersects the rainbow at two points. Choose the points where your drone intersects the parabola and create a table of at least four values for the function. Remember to include the two points of intersection in your table.

OpenStudy (anonymous):

So all you have to do is pick two random numbers and make a "x,y" table

OpenStudy (anonymous):

so how would it be if I pick up -5, -4, -3, -2 and -1?

OpenStudy (anonymous):

How do I know which numbers to use?

OpenStudy (anonymous):

1/4 + w = 6 1/3

OpenStudy (anonymous):

i think 6 1/12

OpenStudy (anonymous):

–2 3/7 + b = 6 1/7

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

I dont think that is the answer for this kind of problem

OpenStudy (anonymous):

pick easy numbers

OpenStudy (anonymous):

\[y=-x^2+36\] pick like \(2,4,6,8\) or sommat

OpenStudy (anonymous):

then pick some nice easy line

OpenStudy (anonymous):

so, I have to pick numbers related to 36 for x?

OpenStudy (anonymous):

no not at all you can pick anything do you know what \(y=-x^2+36\) looks like ?

OpenStudy (anonymous):

yes, I have a picture of it in my assignment, Idk how to post it here though, do you need it?

OpenStudy (anonymous):

no dear i know what it looks like http://www.wolframalpha.com/input/?i=-x^2%2B36+domain+-6..6

OpenStudy (anonymous):

oh i see now we have to make a line that intersects it twice did you pick a line?

OpenStudy (anonymous):

no, I am not sure how to yet, this is very confusing.

OpenStudy (anonymous):